the yearbook club had a meeting. The meeting had 15 people, which is 1/3 of the club. how many people are in the club?
Answer by gwendolyn(128)

We can turn this question into an equation, then solve for the unknown value.
We know that 15 (the number of people who showed up for the meeting) is equal to 1/3 of the value we want to know (the total number of club members)
So, the equation where the total number of members equals x looks like this:
1/3 * x = 15
Now, we just have to divide both sides by 1/3 to get our answer:
x = 15 / (1/3)
Dividing by 1/3 is equal to multiplying by 3, so:
x = 15 * 3
x = 45
Therefore, There are 45 total members in the yearbook club.